A new type of light bulb has an average lifespan of 1200 hours with a standard deviation of 75 hours. Use the table to find the probability that a light bulb chosen at random will last less than 1170 hours.
How do I find probability using a table? What is the formula?

Answer by stanbon(75887)

A new type of light bulb has an average lifespan of 1200 hours with a standard deviation of 75 hours.
Use the table to find the probability that a light bulb chosen at random will last less than 1170 hours.
How do I find probability using a table? What is the formula?
-------
Find the z-value of 1170
z(1170) = (1170-1200)/75 = -0.4
Note: That means 1170 is 0.4 standard deviations below the mean.
-------------------------------------
Your Problem:
P(x < 1170) = P(z < -0.4) = ?
Note: At this point you have to use your z-chart to find the
area under the normal curve to the left of z = -0.4. That
is the same as the area to the right of +0.4.
----
I use a TI-84 calculator to get normalcdf(-100,-0.4) = 0.3446
---------------------
Cheers,
Stan H.